Subject: DIS improvements

From: elt@astro.Princeton.EDU

Submitted: Mon, 27 Jul 1998 16:15:50 -0400 (EDT)

Message number: 62 (previous: 61, next: 63 up: Index)

During the ongoing shutdown, there has been considerable work on 
DIS electronics and mechanics.  I attach below FYI some excerpts
from R. Lucinio's trip report and from a report by Gordon Richards.

Ed Turner


  DIS:  The LN2 fill Forth code was rewritten in order to simplify the time
counter interrupt routine and allow for the generation of code to querry the
system for information on previous fill cycles for the purpose of monitoring
LN2 fill operations and to set or modify the several time constants involved in
the auto-fill system. The new LN2 fill code was run and all fill operations
were verified.
  Note: The Forth command ".LN2", in addition to previous information, now
returns data regarding the previous fill operation, including elapsed time
since the prior fill, conditions triggering the fill sequence, fill times for
each camera, etc. In addition, there is now a new word, "*LN2" which returns
all of the data given by ".LN2" but without the verbiage, just a series of
ASCII values for computer digestion.
   Fiber optic converters for camera control and data aquisition were designed,
built and installed on the DIS Mirella PC camera interface board. Proper
operation was verified and circuit changes were documented, Jon Brinkman
receiving copies of all new DIS Forth code and circuit schematics.

Addendum 24 July 98

  Other DIS problems addressed were :
  Following a report of data aqusition problems, Jon Brinkman and I examened
the test data and saw two problems, one affecting the red camera and one
affecting both cameras.  We felt that the red camera problem was most likely
due to a hot bit probably in the A to D converter. I changed the converter
and Jon ran several test frames which seemed to be free of that problem.
Upon 'scoping the serial clock and data lines, common to both cameras, we
discovered extranious data bits being picked up on the serial data line.
Switching to the last available MAI fiber converter cleared up the symptoms.
Subsequently the MAI converters were abondoned for the on board converters
installed as described above. Data tests run look proper with none of the prior
  There were also a couple of robotics problems involving the grating turret
and its detent. the detent out limit switch needed an adjustment and I re-
defined two of the turret positions, Mirrors and Gratings2 by 1 and 2 steps
respectively.  The robotics were exercised in three orientations of the
instrument and all operations performed without error.


I just got done looking at the bias frames that John Brinkman took
with DIS using the new A/D converter.  If you want to look at the
plot, it is in /home/visitor1/gordonr/dis/

The RN on the red side is now down to about 7.6 ADU (14.9 electrons),
though it would be nice to get a few more bias frames to get a better
measurement of this.  So, congratulations on the fix.  Thanks!

On a similar note, I showed Connie the plots and she noted that the
plot of the blue histogram might be showing that plateau as a result
of the 16 bit flapping around (since the spread is about 16 counts).
Thus it is quite likely that the blue A/D converter needs to be
replaced as well, since there may be a similar problem (albeit a
smaller one) with the A/D converter on the blue side.


APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 62 in the apo35-dis archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to